OpenStudy (anonymous):

If the parent function f(x) = x2 is modified to g(x) = 2x2 + 1, which statement is true about g(x)? It is an even function. It is an odd function. It is both an even and an odd function. It is neither an even nor an odd function.

OpenStudy (anonymous):

f(x)=x^2 is modified to g(x)=2x^2+1

OpenStudy (lexi724):

A function f(x) is an even function if f(-x) = f(x). In other words, replacing x with -x doesn't modify the function. A function f(x) is an odd function if f(-x) = -f(x). In other words, replacing x with -x is identical to the equation obtained after multiplying each term by -1.

OpenStudy (anonymous):

I dont see no -x

OpenStudy (anonymous):

...

OpenStudy (lexi724):

its basically the proof tho

OpenStudy (anonymous):

So is it neither or not

OpenStudy (lexi724):

maybe @iGreen can be of some assistance im terrible at explaining things

OpenStudy (anonymous):

I think its even cause it stretches out

OpenStudy (anonymous):

It has to be even green

OpenStudy (anonymous):

and it is even

OpenStudy (igreen):

I don't know..I'm not sure.

OpenStudy (anonymous):

well it is even

OpenStudy (anonymous):

the answer is even lol